The following problem applies to a perfectly competitive producer of widgets. A typical producer, say Widget Enterpri...

  1. The following problem applies to a perfectly competitive producer of widgets. A typical producer, say Widget Enterprises Inc., can sell widgets at a constant price of $30/pound. Widget Enterprises has the following costs in the short-run. Its total fixed costs are $45.

Quantity Total Cost

(pounds) $

0                                                          45

1                                                          65

2                                                          80

3                                                          90

4                                                        105

5                                                        125

6                                                        150

7                                                        180

8                                                        215

9                                                        255

a.   What does it mean to say that Widget Enterprises is a price taker? What does it say about the widgets it makes and the widgets of other firms?

b. Complete the following schedule:

    Q              TC         TFC    TVC      MC       TR      MR    

(pounds)        $              $           $           $          $         $

0 45

1 65

2 80

3 90

4              105

5              125

6 150

7 180

8 215

9 255

c. Profit is maximized where MR = MC. What is the profit maximizing quantity of widgets?

d. For the answer you gave in part c, compute the profit earned by the firm.

e.   Complete the following schedule:

                                   Q            AFC        AVC         ATC

                             (pounds)    ($/lb)      ($/lb)          ($/lb)         

                                  0              ----         ----             ----

                                  1

                                  2

                                  3

                                  4

                                  5

                                  6

                                  7

                                  8

                                  9

g.   What would the profit maximizing quantity of output be if the price of widgets were $35/lb? What if the price were $40/lb?
Answer #1

Ans) a) Perfectly competitive market are those where there are large number of buyers and sellers. None is big enough to influence the price and hence the price is decided by market forces i.e supply and demand. This is the reason we say that firms in competitive market are price takers.

Also every seller in the market is selling homogeneous products. So here, the widgets sold by all the firms is same.

b) TC= FC+VC

TFC= FC÷Q

TVC= TC-TVC

TVC=TC- TFC

MC= change in total cost÷change in quantity

TR = P×Q

AFC=FC÷Q

AVC=VC÷Q

ATC=TC÷Q

In Perfectly competitive market, marginal revenue (MR) = Price

MC TR AFC AVC ATC TC 45 45 О 65 45 20 1 80 3 45 46 05 65 40 30 20 175 15 15 16 17:5 19.28 6 21-26 23.33 2.33 45 20 30 4 35 15cc) In the fig. Above, we see that at Q=7, MR= MC. So profit maximising quantity is 7.

d) Profit = TR-TC = 210-180 = $30

g)If P= 35, profit maximising quantity is 8 and if P= 40, profit maximising quantity is 9. As at profit maximising point, MR= MC